# Welfare and Cost Aggregation for Multi-Agent Control: When to Choose Which Social Cost Function, and Why?

### v2

@@ IV-C Energy Curtailment
The aforementioned comparability levels dictate which SCF is appropriate to use. (CUC) leaves no other choice than the utilitarian approach of maximizing $\sum_{i}x_{i}$ (which corresponds to OPF-total in [70]). Under (OLC), the appropriate SCF is necessarily $\max_{i}J_{i}(x)$. When $J_{i}(x)=x_{i}$, we recover the Egalitarian curtailment policy proposed in [68]. When $J_{i}(x)=x_{i}/p_{i}$, we recover the OPF-generation policy in [70] (also similar to the Proportional policy in [68]). Finally, when $J_{i}(x)=x_{i}-p_{i}$, we recover the Uniform Dynamic policy in [68] and the OPF-export policy in [70].
@@
Different comparability assumptions justify different policies that have been proposed in the literature, but have not been adequately motivated. Explainability is crucial in this application, as the financial consequences are significant (e.g., the utilitarian approach results in a “water filling” solution where generators are completely curtailed starting from those connected further away from the main grid). Figure 6 provides some intuition on the different curtailment policies. We refer to [68] for a comparison of the numerical solutions to the resulting optimization problems on a benchmark distribution grid.
Figure 6: The effect of different curtailment policies (justified by different comparability assumptions) on five generators. Curtailing generators 4 and 5 is the most efficient way to satisfy the grid operational constraints. Figure adapted from [68].
@@
Inequitable and opaque access to the electricity grid exacerbate existing social disparities [73], hinders investments [74], and may not meet the legal mandate of DSO as transparent and neutral players [75, 76].
@@
Besides providing a principled motivation for different curtailment policies, the proposed approach warns against using fairness indices like those reviewed in [66, 69] (e.g., Gini index, Jain index, F-fairness, K-fairness). While they are useful to measure inequality, they may have unintended consequences when used as social cost functions, as they easily violate basic axioms like (P), i.e., Pareto optimality of the solution.
V Conclusions
@@
Much of applied control work has implicitly been performed with the interests of the population in mind and based on “welfarist” principles as we laid out in this paper. This work aims to provide guidance for making the foundations of the use of social cost functions explicit and, thus, to enable control theorists to express and defend their objectives based on the limits/possibilities of interpersonal comparability in various contexts and applications. In some cases, some social cost functions have to be ruled out, while many functions are permissible in other situations, and our theory thus may restrict or make explicit the range of control objectives that an engineer may pursue. The theoretical and empirical foundations that render different social cost functions preferable (e.g., the efficiency-equity tradeoff in the choice of a SCF under CFC) deserve further investigation.
Appendix A Nash welfare and proportional allocation
Proportional allocation rules (and their relation to proportional fairness and to Nash welfare) have been derived and discussed originally for networking applications, see [58]. In the following, we give a short proof for the special case that is relevant for the water allocation example presented in Section IV-A.
Consider the problem of allocating $\bar{X}$ water resources with each agent having cost $J_{i}(x)=-q_{i}x^{(i)}$ and water rights $c_{i}$. To derive the optimal strategy under Nash social welfare we begin by restating the Nash welfare function as a SCF
$C(\mathbf{J}(x))=-\prod_{i\in\mathcal{N}}\Bigl[-J_{i}(x)\Bigr]^{c_{i}}=-\sum_{i\in\mathcal{N}}c_{i}\log(-J(x))$
under which the water allocation problem becomes
$\min_{x}\;C(\mathbf{J}(x))\quad\text{s.t.}\quad\sum_{i\in\mathcal{N}}x^{(i)}=\bar{X},\;x^{(i)}\geq 0.$
We know that the positive allocation constraint is not active and consequently only dualize the coupling constraint. The corresponding Lagrangian is $L(x,\lambda)=-\sum c_{i}\log(q_{i}x^{(i)})+\lambda(\sum x^{(i)}-\bar{X})$ and by the principle of optimality we have:
$\frac{\partial L}{\partial x^{(i)}}=-\frac{c_{i}}{x^{(i)}}+\lambda=0,$
thus $x^{(i)*}=\frac{c_{i}}{\lambda}$, further primal feasibility yields
$\frac{\sum_{i\in\mathcal{N}}c_{i}}{\lambda}=\bar{X}\;\Rightarrow\;\lambda=\frac{\sum_{i\in\mathcal{N}}c_{i}}{\bar{X}}$
and, consequently, the optimal allocation strategy corresponds to a proportional allocation
$x^{(i)*}=\frac{c_{i}}{\sum_{j\in\mathcal{N}}c_{j}}\,\bar{X},$
where each agent receives a fraction of the total water proportional to their water rights.

### v3

@@ 4.C Energy Curtailment
The aforementioned comparability levels dictate which SCF is appropriate to use. (CUC) leaves no other choice than the utilitarian approach of maximizing $\sum_{i}x_{i}$ (which corresponds to OPF-total in [liu2020fairness]). Under (OLC), the appropriate SCF is necessarily $\max_{i}J_{i}(x)$. When $J_{i}(x)=x_{i}$, we recover the Egalitarian curtailment policy proposed in [gebbran2021fair]. When $J_{i}(x)=x_{i}/p_{i}$, we recover the OPF-generation policy in [liu2020fairness] (also similar to the Proportional policy in [gebbran2021fair]). Finally, when $J_{i}(x)=x_{i}-p_{i}$, we recover the Uniform Dynamic policy in [gebbran2021fair] and the OPF-export policy in [liu2020fairness].
@@
Different comparability assumptions justify different policies that have been proposed in the literature, but have not been adequately motivated. Explainability is crucial in this application, as the financial consequences are significant (e.g., the utilitarian approach results in a “water filling” solution where generators are completely curtailed starting from those connected further away from the main grid). Figure 6 provides some intuition on the different curtailment policies. We refer to [gebbran2021fair] for a comparison of the numerical solutions to the resulting optimization problems on a benchmark distribution grid.
@@
Inequitable and opaque access to the electricity grid exacerbate existing social disparities [Brockway2021], hinders investments [Cuenca2023], and may not meet the legal mandate of DSO as transparent and neutral players [EU2019_944, CERRE2022_DSO].
Figure 6: The effect of different curtailment policies (justified by different comparability assumptions) on five generators. Curtailing generators 4 and 5 is the most efficient way to satisfy the grid operational constraints. Figure adapted from [gebbran2021fair].
@@
Besides providing a principled motivation for different curtailment policies, the proposed approach warns against using fairness indices like those reviewed in [soares2024review, lusis2019reducing] (e.g., Gini index, Jain index, F-fairness, K-fairness). While they are useful to measure inequality, they may have unintended consequences when used as social cost functions, as they violate basic axioms like Pareto optimality of the solution.
5 Conclusions and Outlook
@@
Much of applied control work has implicitly been performed with the interests of the population in mind and based on “welfarist” principles as we laid out in this paper. We aimed to provide guidance for making the foundations of the use of social cost functions explicit and, thus, to enable control theorists to express and defend their objectives based on the limits/possibilities of interpersonal comparability in various contexts and applications. In some cases, some social cost functions have to be ruled out, while many functions are permissible in other situations, and our theory thus may restrict or make explicit the range of control objectives that an engineer may pursue. In practice, the designer should first state the admissible level of interpersonal comparability, then select a social cost function from the class that respects this choice, and report both as part of the problem specification. This turns an implicit convention into an explicit specification and allows one to understand and examine the fairness and efficiency trade-offs.
Multiple questions remain open and should be investigated to enable an effective use of this social-choice-theory perspective in control design problems. We list three directions.
  • Closed-loop certificates – The examples presented in Section 4 are static decision problems. In a control problem, one has to design a feedback policy and then analyze its behavior in the closed-loop interconnection with the plant, taking into account stochastic disturbances, parametric uncertainty, and measurement noise. Closed-loop certificates for the transient and tracking performance need to be derived, employing operations like those proposed in Section 3.C.
  • Online SCF optimization – In dynamic settings, exogenous information is incrementally revealed to the decision maker, agents’ preferences are time-varying, and present decisions affect future trajectories. Social choice results need to be combined with computationally tractable dynamic programming tools to design principled online policies [Kulkarni2020ChangingPreferences, Freeman2017DynamicSettings].
  • Intermediate comparability notions – In realistic settings, the level of comparability between agents’ costs may be revealed through noisy measurements, prior beliefs, and users’ feedback. The standard classification offered by social choice theory is too coarse to capture these settings, and the theory needs to be developed to allow principled social decisions in the face of more nuanced comparability levels.